OpenStudy (anonymous):

determine whether the expression is a polynomial.If it is,state how many terms and variables the polynomial contains.Then state its degree. -6x-1/2x

OpenStudy (anonymous):

im lost pl help me

OpenStudy (anonymous):

Is that 1/(2x) , that is : Is the x in the denominator? Or is that (1/2)x ? It makes a huge difference as one is a polynomial and the other is not.

OpenStudy (anonymous):

Is it \[-6x - \frac{ 1 }{ 2x }\] or \[-6x - \frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/2x

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

That didn't clarify anything 1/2x .... the way your wrote that makes it UNCLEAR whether the x is in the denominator(like in the 1st equation I wrote) or not in the denominator (like the 2nd equation I wrote)

OpenStudy (anonymous):

the down one

OpenStudy (anonymous):

This one? \[-6x - \frac{ 1 }{ 2 }x\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Oh, ok. That is NOT a polynomial. A polynomial can only have terms where the exponent of x is a positive integer. X cannot be in the denominator. These are all polynomials: 4x^2 - 2x + 8, 18x^5 - 9 These are NOT polynomial: 5/x , 3x^(-2) -2x + 5

OpenStudy (anonymous):

the polynomial has __term and __variable